The value of x in the equation log (x-1)9=2 is

Answer by stanbon(48569)

log (x-1)9=2
In exponential form this can be written as follows:
(x-1)^2 = 9
(x-1)^2-9 = 0
Factor to get:
(x-1)-3 = 0 or (x-1)+3=0
x=4 = x=-2
But x cannot be -2 because the base of the log must
be greater than zero.
So, only solution is x=4
--------------
Cheers,
Stan H.